# LLoCO: Learning Long Contexts Offline

### v1

@@ Abstract
Processing long contexts remains a challenge for large language models (LLMs) due to the quadratic computational and memory overhead of the self-attention mechanism and the substantial KV cache sizes during generation. We propose a novel approach to address this problem by learning contexts offline through context compression and in-domain parameter-efficient finetuning. Our method enables an LLM to create a concise representation of the original context and efficiently retrieve relevant information to answer questions accurately. We introduce LLoCO, a technique that combines context compression, retrieval, and parameter-efficient finetuning using LoRA. Our approach extends the effective context window of a 4k token LLaMA2-7B model to handle up to 128k tokens. We evaluate our approach on several long-context question-answering datasets, demonstrating that LLoCO significantly outperforms in-context learning while using 30×30$\times$30 × fewer tokens during inference. LLoCO achieves up to 7.62×7.62$\times$7.62 × speed-up and substantially reduces the cost of long document question answering, making it a promising solution for efficient long context processing. Our code is publicly available at https://github.com/jeffreysijuntan/lloco.
@@ 1 Introduction
Despite the impressive progress made by the LLM model providers, scaling these models to adeptly manage extended contexts remains a formidable challenge, both technically and financially. Due to the self-attention mechanism, transformer-based LLMs incur a quadratic computational and memory overhead as sequence length increases. Many long-context tasks require repeatedly processing over the same prompt, which incurs extra latency overhead and substantial costs, as most commercial LLMs operate on a pricing model that is directly tied to the number of tokens processed. For example, a single inference run with a document that has 100k tokens would take 1.5 USD on Claude 3 Opus [^1] https://www.anthropic.com/api and 1 USD on GPT-4-turbo [^2] https://openai.com/pricing.
Orthogonal to the exiting efforts of expanding the context window limit, our study introduces an innovative strategy to tackle the long context challenge. We propose a method where context information is condensed offline through finetuning, enabling the model to provide accurate responses during inference with streamlined contextual representations. To illustrate our idea, consider an analogy: envision an LLM as a student preparing for an exam, where we, the researchers, are the examiners providing study materials and questions. Traditional in-context learning with full context or Retrieval-Augmented Generation (RAG) resembles an open-book exam, where the LLM has access to all materials while answering questions. In contrast, our approach is akin to a semi-closed-book exam, where the LLM cannot bring the entire book but is allowed to bring a cheat sheet. To excel in the exam, the student must 1) study efficiently to distill a concise yet informative cheat sheet, and 2) effectively retrieve relevant information from the cheat sheet to accurately answer exam questions. Namely, 1) How can we train a model to produce a compact representation of the original context that the LLM can interpret and utilize effectively? 2) How to enable the LLM to proficiently navigate and extract pertinent details from this representation during inference?
Regarding the first question, we find the formulation closely resembles the existing research direction on context compression. Prior works (Chevalier et al., 2023; Mu et al., 2023; Ge et al., 2023) have proposed compressors that aim to distill the essence of the original texts into compact representations that are aligned with LLMs. Our work primarily focuses on the second question – we’ve observed that despite the progress in context compression, LLMs often struggle to accurately read such “cheat sheets” and tend to hallucinate when applying them to answer queries. To address this issue, we employ in-domain parameter-efficient finetuning directly on the compacted context (cheat sheet) without altering its content, which significantly improves the LLM’s ability to accurately extract and utilize information from these compressed representations. In this way, we can steer the LLM to process long context more efficiently and accurately. This approach allows us to extend the effective context window of a 4k LLaMA2-7B model to handle up to 128k tokens. Moreover, we achieve state-of-the-art results that match or even surpass the performance of a LLaMA2-7B-32k model with full context on long context benchmarks, while using 30 times fewer tokens.
@@
Figure 1: The architecture of regular LLM (left) vs LLoCO (right). In regular LLM, the long context gets appended to the prompt directly. In LLoCO, the long context goes through the context encoder first. The output summary token embeddings, which get prepended to the LLM’s prompt, are significantly shorter than the original token embeddings.
@@
This insight has led us to introduce LLoCO, a pipeline that learns contexts offline through the combination of context compression and parameter-efficient finetuning. Our pipeline consists of three stages: preprocessing, finetuning, and serving. First, we preprocess the documents into “cheat sheets”. Then, we employ Low-Rank Adaptation (LoRA) (Hu et al., 2022) to perform parameter-efficient finetuning on these “cheat sheets” in groups. During model serving, we use a standard RAG retriever to retrieve the compressed document as well as the most relevant LoRA module, and apply them to the LLM for inference. The contributions of our work could be summarized as follows:
@@
  • We investigate how to learn long contexts effectively and demonstrate that through a novel combination of context compression and instruction finetuning, we can extend the context window of a 4k LLaMA2-7B to handle context lengths of up to 128k tokens, while achieving performance that greatly surpasses in-context learning with 30 times fewer tokens and 8.34×8.34$\times$8.34 × latency speed-up.
@@
  • We propose LLoCO, a novel pipeline that combines context compression, retrieval, and parameter-efficient finetuning. This pipeline could be deployed to significantly speed up and reduce the cost of long document question answering.
@@ 2 Related work
Long-context LLMs Recently, there have been efforts to increase the LLMs’ context window sizes efficiently with continual pretraining or finetuning. One line of work focuses on scaling the Rotary Position Embeddings (RoPE) (Su et al., 2021), achieving longer contexts up to $128$(Chen et al., 2023b; 2024; Peng et al., 2024). Mistral (Jiang et al., 2023a) proposes sliding window attention that only attends to part of tokens from the previous layer, reducing compute and enabling pretraining with long-context to $30$k. Nevertheless, as the auto-regressive generation in LLMs is largely memory-bound (Kwon et al., 2023), storing the KV cache of longer contexts slows down the inference and requires large GPU VRAMs.
@@
Context compression A closely related topic is context compression, which aims to train a general compressor that can compress any input prompts. GIST (Mu et al., 2023), AutoCompressor (Chevalier et al., 2023), and ICAE (Ge et al., 2023) finetune LLMs in a “soft prompt-tuning” manner, either applying specific regularization in attention masks or utilizing dedicated “memory tokens” to compress contexts into embeddings with significant shorter lengths. LLMLingua family (Jiang et al., 2023c; b; Pan et al., 2024) proposes a question-aware framework to compress prompts in natural languages for black-box LLM APIs. Another line of work employs KV cache compression by eviction (Zhang et al., 2024b; Xiao et al., 2024), which only keeps informative keys and values for generation during inference, or quantization (Sheng et al., 2023b; Hooper et al., 2024). However, all those approaches aim to compress any inputs, which usually incurs rapid performance drops as the compression ratio exceeds a certain limit (e.g. $2$-4×4$\times$4 ×), especially for out-of-distribution texts, hindering their practical applications. In this work, we mainly focus on extreme compression of in-distribution documents of interest, going up to 30×30$\times$30 × compression rates.
@@
Retrieval-augmented Generation Retrieval enhances the performance of LLMs in knowledge-intensive tasks such as question answering with long documents or in open-domain. RAG techniques are effective both in finetuned scenarios and when applied to off-the-shelf LLMs (Guu et al., 2020; Lewis et al., 2020; Zhang et al., 2024a), and have seen many advancements recently, emphasizing improved external knowledge integration, broader query support, and enhanced content clarity (Jiang et al., 2023d; Schick et al., 2023; Asai et al., 2023). Despite the improved performance from retrieval, challenges still remain in terms of runtime efficiency (Mallen et al., 2022) and effective filtering of irrelevant parts (Shi et al., 2023; Xu et al., 2024). Our proposed method opens up new opportunities for capturing important information while ensuring efficient generation.
@@
Parameter-efficient Finetuning Parameter-efficient finetuning methods (Hu et al., 2022; Lester et al., 2021; Liu et al., 2024b) freeze the majority of the model weights and only update a small subset to finetune large models efficiently. LoRA (Hu et al., 2022) is one of the most widely adopted techniques, and recent advancements (Sheng et al., 2023a; Chen et al., 2023a) have focused on enhancing the system performance for deploying LoRA adaptors. In particular, Sheng et al. (2023a) has achieved the deployment of thousands of LoRA instances on a single Nvidia A100 GPU. LLoCO, which utilizes LoRA for finetuning, could benefit from those improvements to deploy LoRA adaptors efficiently in long context scenarios.
3 LLoCO
@@ 3.1 Architecture Overview
Figure 1 illustrates our proposed architecture, which consists of two components: a context encoder and an LLM decoder. In a typical instruction-finetuned LLM, the prompts can be categorized into system, user, and assistant prompts. The system prompt contains task instructions and rules the LLM should follow, as well as relevant contexts. The user prompt is a question asked by the user, and the assistant prompt is the answer generated by the LLM. The context information in the system prompt can include the user’s interaction history or documents related to the user’s question. These contexts can become very long, potentially surpassing the LLM’s context window limit.
@@ 3.2 Pipeline for Offline Context Learning
The pipeline of LLoCO consists of three primary stages: the preprocessing stage, the finetuning stage, and the serving stage. In this section, we introduce the pipeline and how to combine it with an existing RAG system to enable retrieval-augmented document QA with compressed context.
Preprocessing The preprocessing step in a traditional RAG system involves building a vector database to index a collection of documents. This process typically includes three main steps: 1) chunking the documents into passages if they exceed a certain length, 2) generating sentence embeddings of the passages using a text embedding model, and 3) indexing the passages using these embeddings. The result of the preprocessing step is a key-value map that associates each passage with its corresponding embedding.
@@
In LLoCO, we introduce an additional preprocessing step that leverages our context encoder to process the original documents. Since we use AutoCompressor as our context encoder, we follow its practice of dividing the documents into chunks of 1536 tokens and passing them through the context encoder. The context encoder outputs 50 summary embeddings for each chunk recursively, with each embedding having a dimension of 4096. These summary token embeddings are stored alongside the original passages in the vector database, indexed by the sentence embedding of each passage.
@@
Finetuning During the finetuning stage, we first segment the documents into groups based on their type (e.g., academic papers, news) or the tasks that users want to perform on them (e.g., QA, summarization). For each group of documents, we perform parameter-efficient finetuning using a LoRA adaptor. The finetuning data can be in-domain instruction pairs provided by the model provider. If such a dataset does not exist, it could also be generated using self-instruct (Wang et al., 2022) techniques or distilled from a more powerful model like GPT-4. At the end of the finetuning process, we obtain a finetuned LoRA adaptor for each group of documents. In the vector database, we store an identifier for the corresponding LoRA module alongside each passage entry. Additionally, we create a separate LoRA database to store all the LoRA adaptors.
Figure 2: The serving stage of LLoCO’s pipeline.
In Figure 2, we illustrate LLoCO’s serving pipeline. Instead of retrieving the actual passages, we use the retriever to retrieve the compressed token embeddings of the passages. These token-embeddings are concatenated and prepended to the decoder LLM. Additionally, the system searches for the corresponding LoRA adaptor in the database and applies it to the decoder LLM. Applying a LoRA adaptor incurs minimal overhead to the system cost  (Hu et al., 2022). By leveraging recent work on serving multiple LoRA adaptors (Sheng et al., 2023a; Chen et al., 2023a), we can parallelize LoRA adaptor serving and simultaneously serve requests for documents associated with thousands of LoRA adaptors on a single GPU.
In our current system, we assume that all the passages retrieved from the vector database for a given user query belong to the same document group, allowing us to use a single dedicated LoRA adaptor. However, our approach can be extended to handle cases where the retrieved passages span multiple document groups. For instance, it is feasible to design algorithms to dynamically select the most relevant LoRA adaptor based on the majority of the retrieved documents or weight the contributions of different LoRA adaptors according to their relevance to the query (Muqeeth et al., 2024). Another interesting and orthogonal research direction is to explore composing multiple LoRA modules together (Huang et al., 2023), and we leave it as future work.
@@ 4 Experiments
In the experiment section, we aim to investigate the following aspects of LLoCO: (1) its effectiveness in comprehending compressed long contexts, (2) the extent to which summary embeddings can preserve essential information, and (3) the associated system costs.
Datasets We select four datasets dedicated to question-answering tasks—QuALITY (Pang et al., 2021), Qasper (Dasigi et al., 2021), NarrativeQA (Kočiskỳ et al., 2018), and HotpotQA (Yang et al., 2018)—along with one for summarization, QMSum (Zhong et al., 2021). All datasets have long documents as contexts. For all the datasets, we use their validation set for evaluation. We follow the official metrics for each dataset. For QuALITY, we report the exact match (EM) score. For QMSum, we report the geometric mean of ROUGE scores. For the remaining datasets (Qasper, NarrativeQA, and HotpotQA), we report the F1 scores. More details on these datasets can be found in Appendix A.2
@@
4.1 Long Document QA
  1. 1.
    LLaMA-2-7B-4k/32k without Context. In this setting, we do not provide any documents to the LLM in the system prompt.
  2. 2.
    LLaMA-2-7B-4k/32k with Context. In this setting, we provide the LLMs with the ground truth document. We truncate the documents if their length exceeds the context window limit. This setting serves as the baseline.
  3. 3.
  4. 4.
    AutoCompressor. In this setting, we use AutoCompressor (Chevalier et al., 2023) to compress the contexts into summary embeddings and prepend them to the LLM, without performing any in-domain finetuning. This is equivalent to using an AutoCompressor straight out of the box. The AutoCompressor compresses a chunk of 1536 tokens into 50 summary tokens, resulting in an effective context window of roughly 128k tokens. We do not truncate the context unless it exceeds this limit.
  5. 5.
Our results are summarized in Table 1. When AutoCompressor is used out of the box without any in-domain finetuning, its performance sometimes falls short of the baseline where no context is appended. In contrast, by combining compression and finetuning, LLoCO significantly outperforms the baseline on all datasets by a substantial margin while using 30 times fewer tokens. Notably, for NarrativeQA, LLoCO achieves an F1 score of 28.34 compared to 14.42 for the baseline, showcasing the efficacy of our proposed technique.
Furthermore, the effectiveness of LLoCO is particularly evident in the case of the NarrativeQA dataset, where the average document length is 84,770 tokens, significantly exceeding the context window limit of LLaMA2-7B. By compressing the contexts into approximately 2,600 tokens on average, LLoCO is able to fit the compressed representations within the context window limit. This enables LLoCO to leverage all the available context, resulting in impressive F1 scores on this challenging dataset.
@@ 4.1 Long Document QA
4k w.o. Context 4k 1x 32.16 7.68 12.73 10.85 22.22 17.13
32k w.o. Context 32k 1x 31.74 6.30 12.79 10.61 20.03 16.29
4k w. Context 4k 1x 40.45 16.67 14.62 14.42 32.47 23.44
32k w. Context 32k 1x 38.88 21.72 14.58 16.76 31.58 24.70

### v2

@@ Abstract
Processing long contexts remains a challenge for large language models (LLMs) due to the quadratic computational and memory overhead of the self-attention mechanism and the substantial KV cache sizes during generation. We propose LLoCO, a novel approach to address this problem by learning contexts offline through context compression and in-domain parameter-efficient finetuning with LoRA. Our method enables an LLM to create a concise representation of the original context and efficiently retrieve relevant information to answer questions accurately. Our approach extends the effective context window of a 4k token LLaMA2-7B model to handle up to 128k tokens. We evaluate our approach on several long-context question-answering datasets, demonstrating that LLoCO significantly outperforms in-context learning while using 30×30$\times$30 × fewer tokens during inference. LLoCO achieves up to 7.62×7.62$\times$7.62 × speed-up during inference and 11.52×11.52$\times$11.52 × higher throughput during finetuning, substantially reduces the cost of long document question answering. This makes it a promising solution for efficient long context processing. [^1] Our code is publicly available on https://github.com/jeffreysijuntan/lloco.
LLoCO: Learning Long Contexts Offline
Sijun Tan*, Xiuyu Li*, Shishir Patil, Ziyang Wu, Tianjun Zhang, Kurt Keutzer, Joeseph E. Gonzalez, Raluca Ada Popa UC Berkeley {sijuntan,xiuyu}@berkeley.edu
@@ 1 Introduction
Despite the impressive progress made by the LLM model providers, scaling these models to adeptly manage extended contexts remains a formidable challenge, both technically and financially. Due to the self-attention mechanism, transformer-based LLMs incur a quadratic computational and memory overhead as sequence length increases. Many long-context tasks require reusing the same context many times, which incurs extra latency overhead and substantial costs, as most commercial LLMs operate on a pricing model that is directly tied to the number of tokens processed. For example, a single inference run with a document that has 100k tokens would take 1.5 USD on Claude 3 Opus [^2] https://www.anthropic.com/api and 1 USD on GPT-4-turbo [^3] https://openai.com/pricing.
@@
Figure 1: The architecture of regular LLM (left) vs LLoCO (right). In regular LLMs, long contexts are appended directly to the prompt. In contrast, LLoCO first processes these contexts through a context encoder. The resulting summary token embeddings are then prepended to the LLM’s prompt, which are significantly shorter. LLoCO instruction finetunes on embeddings of targeted document groups using a LoRA module. This aligns the LLM’s embedding space with the summary embeddings while keeping both the LLM and context encoder unchanged.
Orthogonal to the exciting efforts of expanding the context window limit, our study introduces an innovative strategy to tackle the long context challenge. We propose a method where context information is condensed offline through finetuning, enabling the model to provide accurate responses during inference with streamlined contextual representations. To illustrate our idea, consider an analogy: envision an LLM as a student preparing for an exam, where we, the researchers, are the examiners providing study materials and questions. Traditional in-context learning with full context or Retrieval-Augmented Generation (RAG) resembles an open-book exam, where the LLM has access to all materials while answering questions. In contrast, our approach is akin to a semi-closed-book exam, where the LLM cannot bring the entire book but is allowed to bring a cheat sheet. To excel in the exam, the student must 1) study efficiently to distill a concise yet informative cheat sheet, and 2) effectively retrieve relevant information from the cheat sheet to accurately answer exam questions. Namely, 1) How can we train a model to produce a compact representation of the original context that the LLM can interpret and utilize effectively? 2) How to enable the LLM to proficiently navigate and extract pertinent details from this representation during inference?
Regarding the first question, we find the formulation closely resembles the existing research direction on context compression. Prior works (Chevalier et al., 2023; Mu et al., 2023; Ge et al., 2023; Yen et al., 2024) have proposed compressors that aim to distill the essence of the original texts into compact representations that are aligned with LLMs. Our work primarily focuses on the second question – we’ve observed that despite the progress in context compression, LLMs often struggle to accurately read such “cheat sheets" and tend to hallucinate when applying them to answer queries. To address this issue, we employ in-domain parameter-efficient finetuning directly on the compacted context (cheat sheet) without altering its content, which significantly improves the LLM’s ability to accurately extract and utilize information from these compressed representations. In this way, we can steer the LLM to process long context more efficiently and accurately. This approach allows us to extend the effective context window of a 4k LLaMA2-7B model to handle up to 128k tokens. Moreover, we achieve state-of-the-art results that match or even surpass the performance of a LLaMA2-7B-32k model with full context on long context benchmarks, while using 30×30$\times$30 × fewer tokens.
@@
This insight has led us to introduce LLoCO, a pipeline that learns contexts offline through the combination of context compression and parameter-efficient finetuning. It consists of three stages: preprocessing, finetuning, and serving. First, we preprocess the documents into “cheat sheets”. Then, we employ Low-Rank Adaptation (LoRA) (Hu et al., 2022) to perform parameter-efficient finetuning on these “cheat sheets” in groups. For serving system design, we use a standard RAG retriever to retrieve the compressed document as well as the most relevant LoRA module, and apply them to the LLM for inference. The contributions of our work could be summarized as follows:
@@
  • We introduce a novel method for effectively modeling long contexts by combining context compression with instruction finetuning. Our approach extends the context window of a 4k LLaMA2-7B model to handle up to 128k tokens, achieving performance that significantly surpasses in-context learning while using 30×30$\times$30 × fewer tokens.
@@
  • We propose LLoCO, a novel framework that combines context compression, retrieval, and parameter-efficient finetuning. This pipeline could be deployed to significantly speed up and reduce the cost of long document question answering. With the compressed context during inference and instruction finetuning, we demonstrate that LLoCO achives a 7.62×7.62$\times$7.62 × speedup over inference latency, and a 11.52×11.52$\times$11.52 × higher throughput compared to finetuning over the original context.
@@ 2 Related work
Long-context LLMs Recently, there have been efforts to increase the LLMs’ context window sizes efficiently with continual pretraining or finetuning. One line of work focuses on scaling the Rotary Position Embeddings (RoPE) (Su et al., 2021), achieving longer contexts up to $128$(Chen et al., 2023b, 2024; Peng et al., 2024). Mistral (Jiang et al., 2023a) proposes sliding window attention that only attends to part of tokens from the previous layer, reducing compute and enabling pretraining with long-context to $30$k. Nevertheless, as the auto-regressive generation in LLMs is largely memory-bound (Kwon et al., 2023), storing the KV cache of longer contexts slows down the inference and requires large GPU VRAMs.
@@
Context compression A closely related topic is context compression, which aims to train a general compressor that can compress any input prompts. GIST (Mu et al., 2023), AutoCompressor (Chevalier et al., 2023), and ICAE (Ge et al., 2023) finetune LLMs in a “soft prompt-tuning" manner, either applying specific regularization in attention masks or utilizing dedicated “memory tokens" to compress contexts into embeddings with significant shorter lengths. LLMLingua family (Jiang et al., 2023c, b; Pan et al., 2024) proposes a question-aware framework to compress prompts in natural languages for black-box LLM APIs. Another line of work employs KV cache compression by eviction (Zhang et al., 2024b; Xiao et al., 2024; Li et al., 2024; Tang et al., 2024), which only keeps informative keys and values for generation during inference, or quantization (Sheng et al., 2023b; Liu et al., 2024c; Hooper et al., 2024). However, those previous approaches aim to compress any inputs, which usually incurs rapid performance drops as the compression ratio exceeds a certain limit (e.g. 4×4$\times$4 ×), especially for out-of-distribution texts. Additionally, these methods may require custom kernel implementations to achieve real-time acceleration, limiting their practical use. In contrast, our work focuses on extreme compression of in-distribution documents, achieving up to 30× compression without the need for custom kernels.
@@
Retrieval-augmented Generation Retrieval enhances the performance of LLMs in knowledge-intensive tasks such as question answering with long documents or in open-domain. RAG techniques are effective both in finetuned scenarios and when applied to off-the-shelf LLMs (Guu et al., 2020; Lewis et al., 2020; Zhang et al., 2024a), and have seen many advancements recently, emphasizing improved external knowledge integration, broader query support, and enhanced content clarity (Jiang et al., 2023d; Schick et al., 2023; Asai et al., 2023). Despite the improved performance from retrieval, challenges still remain in terms of runtime efficiency (Mallen et al., 2022) and effective filtering of irrelevant parts (Shi et al., 2023b; Xu et al., 2024). Our proposed method opens up new opportunities for capturing important information while ensuring efficient generation.
@@
Parameter-efficient Finetuning A line of methods (Hu et al., 2022; Lester et al., 2021; Shi et al., 2023a; Liu et al., 2024b) freezes the majority of the model weights and only updates a small subset to finetune large models efficiently. LoRA (Hu et al., 2022) is one of the most widely adopted techniques, and recent advancements (Sheng et al., 2023a; Chen et al., 2023a) have focused on enhancing the system performance for deploying LoRA adaptors. In particular, Sheng et al. (2023a) has achieved the deployment of thousands of LoRA instances on a single NVIDIA A100 GPU. LLoCO, which utilizes LoRA for finetuning, could benefit from those improvements to deploy LoRA adaptors efficiently in long context scenarios.
3 Method
@@ 3.1 Architecture Overview
Figure 1 illustrates our proposed LLoCO architecture, which consists of two components: a context encoder and an LLM decoder. In a typical instruction-finetuned LLM, the prompts can be categorized into system, user, and assistant prompts. The system prompt contains task instructions and rules the LLM should follow, as well as relevant contexts. The user prompt is a question asked by the user, and the assistant prompt is the answer generated by the LLM. The context information in the system prompt can include the user’s interaction history or documents related to the user’s question. These contexts can become very long, potentially surpassing the LLM’s context window limit.
@@ 3.2 Pipeline for Offline Context Learning
The pipeline of LLoCO consists of two primary stages: the preprocessing stage and the finetuning stage. We also outline a serving stage for real-world deployment in the system design.
@@
Preprocessing First, we employ a preprocessing stage that leverages our context encoder to process the original documents. Since AutoCompressor is used as our context encoder, we follow its practice of dividing the documents into chunks of 1536 tokens and passing them through the context encoder. The context encoder outputs 50 summary embeddings for each chunk recursively, with each embedding having a dimension of 4096.
Our preprocessing stage is designed to be extensible, allowing seamless integration into a retrieval-augmented generation (RAG) system for document QA. In a typical RAG system, preprocessing involves building a vector database to index a collection of documents, where documents are chunked into passages, and each passage is associated with a key— a sentence embedding generated by a text embedding model.The design of LLoCO allows summary token embeddings to be stored alongside the original passages in the vector database, indexed by the same passage key.
@@
Finetuning During the finetuning stage, we first segment the documents into groups based on their type (e.g., academic papers, news) or the tasks that users want to perform (e.g., QA, summarization). For each group of documents, we perform parameter-efficient finetuning using a LoRA adaptor. The finetuning data can be in-domain instruction pairs provided by the model provider. If such a dataset does not exist, it could also be generated using self-instruct (Wang et al., 2022) techniques or distilled from a more powerful model like GPT-4. At the end of the finetuning process, we obtain a finetuned LoRA adaptor for each group of documents. In the vector database system design, each passage entry will include an identifier for the corresponding LoRA module. Additionally, a separate database will be established to store all adaptors.
Concretely, given summary tokens $\mathbf{X}_{m}$ and instruction pairs $\{\mathbf{X}_{q}^{1},\mathbf{X}_{a}^{1},\dots,\mathbf{X}_{q}^{L},\mathbf{X}_{a}^{L}\}$ of length $L$ for a document group $g$, we aim to maximize the probability of generating $\mathbf{X}_{a}$ defined as:
$p(\mathbf{X}_{a}|\mathbf{X}_{m},\mathbf{X}_{q})=\prod_{i=1}^{L}p_{\theta_{g}}(\mathbf{X}_{a}^{i}|\mathbf{X}_{m},\mathbf{X}_{q}^{i})$ (1)
where $\theta_{g}$ denotes the LoRA weight for group $g$.
Serving We also design a serving stage that can be naturally extended to a RAG system, where we leverage a standard RAG retriever to retrieve relevant documents, but instead prepend the compressed embeddings of these documents to the LLM decoder. Additionally, we apply the most relevant LoRA module to the model. Readers can refer to Appendix E for more details.
@@ 4 Experiments
In the experiment section, we aim to investigate the following aspects of LLoCO: (1) its effectiveness in comprehending compressed long contexts, (2) the extent to which summary embeddings can preserve essential information, and (3) the efficiency improvements on the associated system costs.
Datasets We select four datasets dedicated to question-answering tasks—QuALITY (Pang et al., 2021), Qasper (Dasigi et al., 2021), NarrativeQA (Kočiskỳ et al., 2018), and HotpotQA (Yang et al., 2018)—along with one for summarization, QMSum (Zhong et al., 2021). All datasets have long documents as contexts. For all the datasets, we use their validation set for evaluation. We follow the official metrics for each dataset. For QuALITY, we report the exact match (EM) score. For QMSum, we report the geometric mean of ROUGE scores. For the remaining datasets (Qasper, NarrativeQA, and HotpotQA), we report the F1 scores. More details on these datasets can be found in Appendix B.
@@
4k w. Original Context 4k 1x 40.45 16.67 14.62 14.42 32.47 23.44
32k w. Original Context 32k 1x 38.88 21.72 14.58 16.76 31.58 24.70
128k w. Original Context 128k 1x 33.89 20.38 13.88 28.22 27.69 24.81
@@
4.1 Long Document QA
  1. 1.
    LLaMA-2-7B-4k/32k/128k with Original Context. This is a baseline setting where we provide the LLMs with the ground truth document. We truncate the documents if their length exceeds the context window limit.
  2. 2.
  3. 3.
    AutoCompressor. In this setting, we use AutoCompressor (Chevalier et al., 2023) to compress the contexts into summary embeddings and prepend them to the LLM, without performing any in-domain finetuning, which is equivalent to using it straight out of the box. The AutoCompressor compresses a chunk of 1536 tokens into 50 summary tokens, resulting in an effective context window of roughly 128k tokens. We do not truncate the context unless it exceeds this limit.
  4. 4.
Our results are summarized in Table 1. When AutoCompressor is used without in-domain finetuning, its performance sometimes falls short of the baseline where no context is appended. However, by combining compression and finetuning, LLoCO significantly outperforms the baseline on all datasets, using 30×30$\times$30 × fewer tokens.
For the QuALITY, Qasper, QMSum, and HotpotQA datasets, most samples fit within the 32k token limit of the LLaMA2-7B-32k model without truncation. Here, LLoCO does not have a longer context advantage but still uses 30×30$\times$30 × fewer tokens than the baseline while achieving better performance. This shows that in-domain finetuning enhances the model’s ability to interpret compressed embeddings and answer questions. Detailed sequence length statistics are in Table 6 in Appendix B.
For the NarrativeQA dataset, the average document length is 84,770 tokens, exceeding the LLaMA2-7B-32k context limit. LLoCO compresses these contexts to about 2,600 tokens, enabling comprehensive context utilization for question answering. To assess the impact of context length and ensure fair comparison, we conducted an ablation study with two new configurations: (1) documents truncated to 32k tokens before compression and finetuning LLoCO on the resulting embeddings, and (2) testing another LLaMA2-7B extended for $128$k context window with state-of-the-art long-context ability from (Fu et al., 2024). Our results in Table 2 show that LLoCO performs better than LLaMA-7B at both 32k and 128k context lengths.
Setup NQA
LLaMA2-7B-32k (Li et al., 2023) 16.76
LLoCO-32k 27.99
LLaMA2-7B-128k (Fu et al., 2024) 28.22
LLoCO-128k 28.34
Table 2: Impact of context lengths on NarrativeQA.